OpenStudy (anonymous):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (anonymous):

well the first one has a slope of 5 not -5 so throw that one away then you are left with 3 choices. test then the point (-2, 4) by plug in the values for x and y in the given equations to see which one holds

OpenStudy (anonymous):

B for example you have 5x+y=-6 plug in (-2, 4) i have 5(-2)+4=-6 i ask is this true to check i carry on the arithmetic -20+6=-14 so not true

OpenStudy (anonymous):

So on and so forth...

OpenStudy (anonymous):

only one will hold!

OpenStudy (anonymous):

Thank you! Could you help with this one? A line has a slope of and a y-intercept of 3. Which equations represent the line? Choose all answers that are correct. A. 2y=6+x B. x+2y=6 C. y=-1/2+3 D.(y-2)=-1/2(x-2) I know that A is wrong and B and C are right but I'm not sure about D.